OpenStudy (anonymous):

y'= 2 sqrt (y) sin(x)... can you differentiate again this??

OpenStudy (anonymous):

we must to do with product rule right? ... iam stuck at \[y''=2\sqrt{y}\times \cos(x) + \sin(x) \times \frac{ y' }{ \sqrt{y} }\]

OpenStudy (nincompoop):

second-order implicit differentiation?

OpenStudy (anonymous):

and then? for y' we get.... \[y'=2\sqrt{y} \sin(x)\] ...so just substitutes it into y'' ... and then..i dont know how? can help me?? @ganeshie8 @jhonyy9 @Kbutler40 @kc_kennylau @rvc @radar

OpenStudy (anonymous):

@nincompoop yups... second implicit... can u help me

OpenStudy (kc_kennylau):

What do you get when you substitute?

OpenStudy (anonymous):

iam blur... beause lastly we must to substitutes the answer with the first equation..... first equation is: cos(x)+ sqrt y =5

OpenStudy (kc_kennylau):

What's the original question anyway

OpenStudy (anonymous):

and after that..differentiate... I get y'=2 sqrt(y) sin(x).. after that must to differentiate again... but iam stuck... original question is cos x + sqrt(Y)=5

ganeshie8 (ganeshie8):

y'=2 sqrt(y) sin(x)

OpenStudy (kc_kennylau):

cos x + sqrt(y) = 5 can't be a question, what's the original question

OpenStudy (kc_kennylau):

the full question please

ganeshie8 (ganeshie8):

substitute sqrt(y) value in y'

ganeshie8 (ganeshie8):

y'=2 sqrt(y) sin(x). y' = 2[5-cos(x)]sin(x) = 10 - 2cos(x)sin(x) = 10 - sin(2x)

ganeshie8 (ganeshie8):

differentiate this again, i think u dont want to see y,y' terms left over in y''

OpenStudy (nincompoop):

you only differentiate implicitly once, the second or higher order of differentiation turns out to be just like explicit

rvc (rvc):

 so by using product rule we can differentiate